# Pixel Is Not A Barrier: An Effective Evasion Attack for Pixel-Domain Diffusion Models

### v1

@@ Abstract
Diffusion Models have emerged as powerful generative models for high-quality image synthesis, with many subsequent image editing techniques based on them. However, the ease of text-based image editing introduces significant risks, such as malicious editing for scams or intellectual property infringement. Previous works have attempted to safeguard images from diffusion-based editing by adding imperceptible perturbations. These methods are costly and specifically target prevalent Latent Diffusion Models (LDMs), while Pixel-domain Diffusion Models (PDMs) remain largely unexplored and robust against such attacks. Our work addresses this gap by proposing a novel attacking framework with a feature representation attack loss that exploits vulnerabilities in denoising UNets and a latent optimization strategy to enhance the naturalness of protected images. Extensive experiments demonstrate the effectiveness of our approach in attacking dominant PDM-based editing methods (e.g., SDEdit) while maintaining reasonable protection fidelity and robustness against common defense methods. Additionally, our framework is extensible to LDMs, achieving comparable performance to existing approaches.
@@ 1 Introduction
In recent years, Generative Diffusion Models (GDMs) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2021) emerged as powerful generative models that can produce high-quality images, propelling advancements in image editing and artistic creations. The ease of using these models to edit (Meng et al. 2021; Wang, Zhao, and Xing 2023; Zhang et al. 2023) or synthesize new image (Dhariwal and Nichol 2021; Rombach et al. 2022) has raised concerns about potential malicious usage and intellectual property infringement. For example, malicious users could effortlessly craft fake images with someone’s identity or mimic the style of a specific artist. An effective protection against these threats is regarded as the diffusion model generating corrupted images or unrelated images to original inputs. Researchers have made significant strides in crafting imperceptible adversarial perturbation on images to safeguard them from being edited by diffusion-based models.
@@
Previous works such as PhotoGuard (Salman et al. 2023) and Glaze (Shan et al. 2023) have effectively attacked Latent Diffusion Models (LDMs) by minimizing the latent distance between the protected images and their target counterparts. PhotoGuard first introduce attacking either encoders or diffusion process in LDMs via Projected Gradient Descent (PGD) (Madry et al. 2018) for the protection purpose; however, it requires backpropagating the entire diffusion process, making it prohibitively expensive. Subsequent works AdvDM (Liang et al. 2023) and Mist (Liang and Wu 2023) leverage the semantic loss and textural loss combined with Monte Carlo method to craft adversarial images both effectively and efficiently. Diff-Protect  (Xue et al. 2024a) further improve adversarial effectiveness and optimization speed via Score Distillation Sampling (SDS)  (Poole et al. 2022), setting the state-of-the-art performance on LDMs.
@@
However, previous works primarily focus on LDMs, and attacks on Pixel-domain Diffusion Models (PDMs) remain largely unexplored. Xue et al.  (Xue et al. 2024a) also highlighted a critical limitation of current methods: the attacking effectiveness is mainly attributed to the vulnerability of the VAE encoders in LDM; however, PDMs don’t have such encoders, making current methods hard to transfer to PDMs. The latest work (Xue and Chen 2024) has attempted to attack PDMs, but the result suggests that PDMs are robust to pixel-domain perturbations. Our goal is to mitigate the gap between these limitations.
@@
In this paper, we propose an innovative framework designed to effectively attack PDMs. Our approach includes a novel feature attacking loss that exploits the vulnerabilities in denoising UNet to distract the model from recognizing the correct semantics of the image, a fidelity loss that acts as optimization constraints that ensure the imperceptibility of adversarial image and controls the attack budget, and a latent optimization strategy utilizing victim-model-agnostic VAEs to further enhance the naturalness of our adversarial image closer to the original. With extensive experiments on different PDMs, the results show that our method is effective and affordable while robust to traditional defense methods and exhibiting attack transferability in the black-box setting.
In addition, our approach outperforms current semantic-loss-based and PGD-based methods, reaching state-of-the-art performance on attacking PDMs. Our contributions are summarized as follows:
@@ 2.1 Evasion Attack for Diffusion Model
The prevalence of diffusion-based image editing techniques has posed the challenges of protecting images from maliciously editing. These editing techniques are mainly based on SDEdit (Meng et al. 2021) or its variations which can be applied to both PDM and LDM to produce the editing results. In general, the editing starts from first transforming the clean image (or the clean latent) into the noisy one by introducing Gaussian noise for the forward diffusion process followed by performing a series of reverse diffusion sampling steps with new conditions. In addition, based on different number of forward diffusion steps, the method could control the extent of the editing results obeying the new conditions while preserving the original image semantics. Notably, a small number of forward steps allow the editing results faithful to the original image, and more variations are introduced when larger forward step value is applied.
To counteract SDEdit-based editing, H. Salman et al. first proposed PhotoGuard (Salman et al. 2023) to introduce two attacking paradigms based on Projected Gradient Descent (PGD) (Madry et al. 2018). The first is the Encoder Attack, which aims to disrupt the latent representations of the Variational Autoencoder (VAE) of the LDMs, and the second is the Diffusion Attack, which focuses more on disrupting the entire diffusion process of the LDMs.
The Encoder Attack is simple yet effective, but the attacking results are sub-optimal due to its less flexibility for optimization than the Diffusion Attack. Although the Diffusion Attack achieves better attack results, it is prohibitively expensive due to its requirement of backpropagation through all the diffusion steps. In the following, we further introduce more relevant work for these attacks along with another common attack for diffusion models.
@@
On the other hand, researchers found that VAEs in widely adopted LDMs are more vulnerable to attack at a lower cost while avoiding attacking the expensive diffusion process. (Salman et al. 2023; Liang and Wu 2023; Shan et al. 2023; Xue et al. 2024b), focus on disrupting the latent representation in LDM via PGD and highlights the encoder attacks are more effective against LDMs.
@@
Most of the LDMs contain conditional modules for steering generation, previous works (Shan et al. 2023, 2024; Lo et al. 2024) exploited the vulnerability of text conditioning modules. By disrupting the cross-attention between text concepts and image semantics, these methods could effectively interfere with the diffusion model for capturing the image-text alignment, thereby realizing the attack.
@@
To our knowledge, previous works primarily focus on adversarial attacks for LDMs, while attacks on PDMs remain unexplored. Xue et al. (Xue and Chen 2024) further emphasized the difficulty of attacking PDMs. However, in our work, we find that by crafting an adversarial image to corrupt the intermediate representation of diffusion UNet, we can achieve promising attack performance for PDMs, while the attack is also compatible with LDMs. Moreover, inspired by (Laidlaw, Singla, and Feizi 2021; Liu et al. 2023) which utilizes LPIPS (Zhang et al. 2018) as distortion measure, we also propose a novel attacking loss as the measure to craft better adversarial images for PDMs.
3 Background on Diffusion Models
Score-based models and diffusion models allow to generate samples starting from easy-to-sample Gaussian noise to complex target distributions via iteratively applying the score function of learned distribution during sampling, i.e., the gradient of underlying probability distribution $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ with respect to $\mathbf{x}$. However, the exact estimations of the score functions are intractable.
To bypass this problem, Yang Song et al. proposed slice score matching (Song et al. 2020), and Ho et al. proposed Denoising Diffusion Probability Model (DDPM) (Ho, Jain, and Abbeel 2020) that first gradually perturbs the clean data with linear combinations of Gaussian noise and clean data, as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ via the predefined timestep schedulers where $t\in[0,T]$ and $\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$, then they finally become isotropic Gaussian noise as time reaches $T$, this is also referred as forward diffusion. The goal is to train a time-dependent neural network that can learn to denoise noisy samples given corresponding timestep $t$. Specifically, the training objective is the expectation over noise estimation MSE, which is formulated as $\mathbb{E}_{t,\mathbf{x},\epsilon_{t}}[\|\epsilon_{t}-\epsilon_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}]$, where $\epsilon_{\theta}$ denotes the parametrized neural network, DDPM adopted UNet (Ronneberger, Fischer, and Brox 2015) as their noise estimating network. During inference time, we first generate a random Gaussian sample, then iteratively apply the noise estimation network $\epsilon_{\theta}$ and perform denoising operations to generate a new clean sample of the learned distribution. Particularly, Song et.al proposed DDIM (Song, Meng, and Ermon 2021) that generalized the DDPM sampling formulation as:
$\mathbf{x}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$ (1)
With $\sigma_{t}=\sqrt{(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}\cdot$ $\sqrt{1-\bar{\alpha}_{t}/\bar{\alpha}_{t-1}}$, Equation 1 becomes DDPM, and when $\sigma_{t}=0$, the sampling process become deterministic as proposed in DDIM since the added noise during each sampling step is null.
@@ 4.1 Threat Model and Problem Setting
The malicious user collects an image $\mathbf{x}$ from the internet and uses SDEdit (Meng et al. 2021) to generate unauthorized image translations or editing, denoted as $\text{SDEdit}(\mathbf{x},t)$, that manipulates the original input image $\mathbf{x}$.
Our work aims to safeguard the input image $\mathbf{x}$ from the authorized manipulations by crafting an adversarial image $\mathbf{x}^{\operatornamewithlimits{adv}}$ by adding imperceptible perturbation to disrupt the reverse diffusion process of SDEdit for corrupted editions.
For example, we want the main object of the image, e.g., the cat in the source image $\mathbf{x}$ as shown in Figure 2 unable to be reconstructed by the reverse diffusion process. Meanwhile, the adversarial image should maintain similarity to the source image to ensure fidelity. The reason why we target SDEdit as our threat model is that it is recognized as the most common and general operation in diffusion-based unconditional image translations and conditional image editing. Additionally, it has been incorporated into various editing pipelines (Tsaban and Passos 2023; Zhang et al. 2023). Here we focus on the unconditional image translations for our main study, as they are essential in both unconditional and conditional editing pipelines. Formally, our objective to effectively safeguard images while maintaining fidelity is formulated as:
@@
where $\mathcal{M}$ indicates natural image manifold, $d$, $d^{\prime}$ indicate image distance functions and $\epsilon$ denotes the fidelity budget.
In the following sections, we first present a conceptual illustration of our method, followed by our framework for solving the optimization problem. We then discuss the novel design of our attacking loss and fidelity constraints, which provide more efficient criteria compared to previous methods for solving the image protection optimization problem using PGD. Finally, we introduce an advanced design to enhance image protection quality by latent optimization via victim-model-agnostic VAE.
Figure 3: Overview of our AtkPDM+ algorithm: Starting from the leftmost latent of the initial adversarial image $\mathbf{z}^{\operatornamewithlimits{adv}}$, we first decode back to pixel-domain to perform forward diffusion with both $\mathbf{x}$ and $\mathbf{x}^{\operatornamewithlimits{adv}}$ and feed them to frozen victim UNet. We then extract the feature representation in UNet to calculate our $\mathcal{L}_{\text{attack}}$, aiming to distract the recognition of image semantics. We also calculate our $\mathcal{L}_{\text{fidelity}}$ in pixel-domain to constrain the optimization. Finally, the $\mathbf{z}^{\operatornamewithlimits{adv}}$ is being alternatively updated by loss gradients.
@@ 4.2 Overview
To achieve effective protection against diffusion-based editing, we aim to push the protected sample away from the original clean sample by disrupting the intermediate step in the reverse diffusion process. For practical real-world applications, it’s essential to ensure the protected image is perceptually similar to the original image. In practice, we uniformly sample the value of the forward diffusion step $t\sim[0,T]$ to generate noisy images and then perform optimization to craft the adversarial image $\mathbf{x}^{\operatornamewithlimits{adv}}$ via our attacking and fidelity losses, repeating the same process $n$ times or until convergence. Figure 2 depicts these two push-and-pull criteria during different noise levels, the successful attack is reflected in the light orange line where the reverse sample moves far away from the normal edition of the image. More specifically, our method can be formulated as follows:
$\max_{\mathbf{x}^{\operatornamewithlimits{adv}}\in\mathcal{M}}\mathbb{E}_{t,\mathbf{x}_{t}|\mathbf{x},\mathbf{x}_{t}^{\operatornamewithlimits{adv}}|\mathbf{x}}\left[-\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})\right]$ (3)
@@
Our framework is illustrated in Figure 3. We fix two identical victim UNets to extract feature representations of clean and protected samples for optimizing to push away from each other. A protection loss is jointly incorporated to constrain the optimization. After $N$ iterations, we segment out only the protecting main object of the image for better imperceptibility of image protection.
@@ 4.3 Proposed Losses
We propose two novel losses as optimization objectives to craft an adversarial example efficiently without running through all the diffusion steps. Attacking loss is designed to distract the feature representation in denoising UNet; Protection loss is a constraint to ensure the image quality. For notation simplicity, we first define the samples $\mathbf{x},\mathbf{x}^{\operatornamewithlimits{adv}}$ in different forwarded steps.
Let $\mathcal{F}(\mathbf{x},t,\epsilon)=\sqrt{\bar{\alpha}_{t}}\mathbf{x}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ be the diffusion forward process. Given timestep $t$ sample from $[0,T]$, noises $\epsilon_{1},\epsilon_{2}$ sample from $\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote $\mathbf{x}_{t}=\mathcal{F}(\mathbf{x},t,\epsilon_{1})$, and $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}=\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon_{1})$.
@@
Our goal is to define effective criteria that could finally distract the reverse denoising process. PhotoGuard (Salman et al. 2023) proposed to backpropagate through all the steps of the reverse denoising process via PGD, however, this approach is prohibitively expensive, Diff-Protect (Xue et al. 2024b) proposed to avoid the massive cost by leveraging Score Distillation (Poole et al. 2022) in optimization. However, Diff-protect relies heavily on gradients of attacking encoder of an LDM as stated in their results. In PDM, we don’t have such an encoder to attack; nevertheless, we find that the denoising UNet has a similar structure to encoder-decoder models, and some previous works (Lin and Yang 2024; Li et al. 2023) characterize this property to accelerate and enhance the generation. From our observations of the feature roles in denoising UNets, we hypothesize that distracting specific inherent feature representation in UNet blocks could lead to effectively crafting an adversarial image. In practice, we first extract the feature representations of forwarded images $\mathbf{x}_{t}$ and $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}$ in frozen UNet blocks of timestep $t$. Then, we adopt 2-Wasserstein distance (Arjovsky, Chintala, and Bottou 2017) to measure the discrepancy in feature space. Note that we take the negative of the calculated distance, as we aim to pull the $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}$ away from $\mathbf{x}_{t}$. Formally, the attacking loss $\mathcal{L}_{\text{attack}}$ is defined as:
$\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}^{\operatornamewithlimits{adv}}_{t})=-\mathcal{W}_{2}\left(\mathcal{U}^{\text{(mid)}}_{\theta}(\mathbf{x}_{t}),\mathcal{U}^{\text{(mid)}}_{\theta}(\mathbf{x}^{\operatornamewithlimits{adv}}_{t})\right).$ (4)
Assuming the feature distributions approximate Gaussian distributions expressed by mean $\mu_{t}$ and $\mu_{t}^{\operatornamewithlimits{adv}}$, and non-singular covariance matrices $\Sigma_{t}$ and $\Sigma_{t}^{\operatornamewithlimits{adv}}$. The calculation of the 2-Wasserstein distance between two normal distributions is viable through the closed-form solution (Dowson and Landau 1982; Olkin and Pukelsheim 1982; Chen, Georgiou, and Tannenbaum 2018):
@@
To control the attack budget for adversarial image quality, we design a constraint function that utilizes the feature extractor from a pretrained classifier for calculating fidelity loss. In our case, we sum up the 2-Wasserstein feature losses of $L$ different layers. Specifically, we define $\mathcal{L}_{\text{fidelity}}$ as:

### v3

@@ Abstract
Diffusion Models have emerged as powerful generative models for high-quality image synthesis, with many subsequent image editing techniques based on them. However, the ease of text-based image editing introduces significant risks, such as malicious editing for scams or intellectual property infringement. Previous works have attempted to safeguard images from diffusion-based editing by adding imperceptible perturbations. These methods are costly and specifically target prevalent Latent Diffusion Models (LDMs), while Pixel-domain Diffusion Models (PDMs) remain largely unexplored and robust against such attacks. Our work addresses this gap by proposing a novel attack framework, AtkPDM. AtkPDM is mainly composed of a feature representation attacking loss that exploits vulnerabilities in denoising UNets and a latent optimization strategy to enhance the naturalness of adversarial images. Extensive experiments demonstrate the effectiveness of our approach in attacking dominant PDM-based editing methods (e.g., SDEdit) while maintaining reasonable fidelity and robustness against common defense methods. Additionally, our framework is extensible to LDMs, achieving comparable performance to existing approaches. Our project page is available at https://alexpeng517.github.io/AtkPDM.
@@ 1 Introduction
In recent years, Generative Diffusion Models (GDMs) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2021) emerged as powerful generative models that can produce high-quality images, propelling advancements in image editing and artistic creations. The ease of using these models to edit (Meng et al. 2021; Wang, Zhao, and Xing 2023; Zhang et al. 2023) or synthesize new images (Dhariwal and Nichol 2021; Rombach et al. 2022) has raised concerns about potential malicious usage and intellectual property infringement. For example, malicious users could effortlessly craft fake images with someone’s identity or mimic the style of a specific artist. An effective protection against these threats is to craft an adversarial image to force the diffusion model to generate corrupted images or unrelated images to the original inputs. Researchers have made significant strides in crafting imperceptible adversarial perturbations on images to protect against diffusion-based editing.
@@
Previous works such as PhotoGuard (Salman et al. 2023) and Glaze (Shan et al. 2023) have effectively attacked Latent Diffusion Models (LDMs) by minimizing the latent distance between the protected images and their target counterparts. PhotoGuard first introduces attacking either encoders or diffusion process in LDMs via Projected Gradient Descent (PGD) (Madry et al. 2018) for the protection purpose; however, it requires backpropagating the entire diffusion process, making it prohibitively expensive. Subsequent works AdvDM (Liang et al. 2023) and Mist (Liang and Wu 2023) leverage the semantic loss and textural loss combined with Monte Carlo method to craft adversarial images both effectively and efficiently. Diff-Protect  (Xue et al. 2024) further improve adversarial effectiveness and optimization speed via Score Distillation Sampling (SDS)  (Poole et al. 2023), setting the state-of-the-art performance on LDMs.
@@
However, previous works primarily focus on LDMs, and attacks on Pixel-domain Diffusion Models (PDMs) remain unexplored. Xue et al.  (Xue et al. 2024) also highlighted a critical limitation of current methods: the attacking effectiveness is mainly attributed to the vulnerability of the VAE encoders in LDM; however, PDMs don’t have such encoders, making current methods hard to transfer to PDMs. The latest work (Xue and Chen 2024) has attempted to attack PDMs, but the result suggests that PDMs are robust to pixel-domain perturbations. Our goal is to mitigate the gap between these limitations.
@@
In this paper, we propose an innovative framework AtkPDM, to effectively attack PDMs. Our approach includes a novel feature attacking loss that exploits the vulnerabilities in denoising UNet to distract the model from recognizing the correct semantics of the image, a fidelity loss that acts as optimization constraints that ensure the imperceptibility of adversarial image and controls the attack budget, and a latent optimization strategy utilizing victim-model-agnostic VAEs to further enhance the naturalness of our adversarial image. With extensive experiments on different PDMs, the results show that our method is effective and affordable while robust to prevalent defense methods and exhibiting attack transferability in the black-box setting. In addition, our approach outperforms current semantic-loss-based and PGD-based methods, reaching state-of-the-art performance on attacking PDMs. Our contributions are summarized as follows:
@@ 2 Related Work
2.1 Image Editing with SDEdit-based Methods
With the multi-step sampling nature and the ease of converting a sample to intermediate noisy latent via forward diffusion of Diffusion Models (Ho, Jain, and Abbeel 2020). SDEdit  (Meng et al. 2021) indicates that the diffusion model sampling process is not necessarily required to begin with random Gaussian noise, but allows starting with a mixture of input image and noise at arbitrary strength, i.e. forwarded to $t\in[0,T]$, for the editing. This technique is generalized to both PDMs and LDMs. Subsequent editing frameworks (Hertz et al. 2023; Tumanyan et al. 2023; Parmar et al. 2023; Mokady et al. 2023) also build upon this concept.
@@ 2.2 Evasion Attack for Diffusion Model
To counteract SDEdit-based editing, Salman et al. first proposed PhotoGuard (Salman et al. 2023) to introduce two attacking paradigms based on Projected Gradient Descent (PGD) (Madry et al. 2018). The first is the Encoder Attack, which aims to disrupt the latent representations of the Variational Autoencoder (VAE) of the LDMs, and the second is the Diffusion Attack, which focuses more on disrupting the entire diffusion process of the LDMs. The Encoder Attack is simple yet effective, but the attacking results are sub-optimal due to its less flexibility for optimization than the Diffusion Attack. Although the Diffusion Attack achieves better attack results, it is prohibitively expensive due to its requirement of backpropagation through all the diffusion steps. In the following, we introduce other proposed method targeting different aspects for attacking diffusion models.
@@
On the other hand, researchers found that VAEs in widely adopted LDMs are more vulnerable to attack at a lower cost than the expensive diffusion process. Hence, they (Salman et al. 2023; Liang and Wu 2023; Shan et al. 2023; Xue et al. 2023) focus on disrupting the latent representation in LDM via PGD and highlight the encoder attacks are more effective against LDMs.
@@
Most of the LDMs contain conditional modules for steering generation, previous works (Shan et al. 2023, 2024; Lo et al. 2024) exploited the vulnerability of text conditioning modules. By disrupting the cross-attention between text concepts and image semantics, these methods effectively interfere with the diffusion model’s ability to capture image-text alignment, thereby achieving the attack.
@@
To the best of our knowledge, previous works primarily focus on adversarial attacks for LDMs, while attacks on PDMs remain unexplored. Xue et al. (Xue and Chen 2024) further emphasized the difficulty of attacking PDMs. However, in our work, we find that by crafting an adversarial image to corrupt the intermediate representation of diffusion UNet, we can achieve promising attack performance for PDMs, while the attack is also compatible with LDMs. Moreover, inspired by (Laidlaw, Singla, and Feizi 2021; Liu et al. 2023) which utilize LPIPS (Zhang et al. 2018) as the distortion measure, we also propose a novel attacking loss as the measure to craft better adversarial images for PDMs.
@@ 3.1 Threat Model and Problem Setting
The malicious user collects an image $\mathbf{x}$ from the internet and uses SDEdit (Meng et al. 2021) to generate unauthorized image translations or editing, denoted as $\text{SDEdit}(\mathbf{x},t)$, that manipulates the original input image $\mathbf{x}$. Our work aims to safeguard the input image $\mathbf{x}$ from the unauthorized manipulations by crafting an adversarial image $\mathbf{x}^{\operatornamewithlimits{adv}}$ through adding imperceptible perturbation to disrupt the reverse diffusion process of SDEdit for corrupted editions. For example, we want the main object of the image, e.g., the cat in the source image $\mathbf{x}$ as shown in Figure 2 is unable to be reconstructed by the reverse diffusion process. Meanwhile, the adversarial image should maintain similarity to the source image to ensure fidelity. The reason why we target SDEdit as our threat model is that it is recognized as the most common and general operation in diffusion-based unconditional image translation and conditional image editing. Additionally, it has been incorporated into various editing pipelines (Tsaban and Passos 2023; Zhang et al. 2023). Here we focus on the unconditional image translation for our main study, as they are essential in both unconditional and conditional editing pipelines. Formally, our objective to effectively safeguard images while maintaining fidelity is formulated as:
@@
where $\mathcal{M}$ indicates natural image manifold, $d$ and $d^{\prime}$ indicate image distance functions, and $\delta$ denotes the fidelity budget.
In the following sections, we first present a conceptual illustration of our method, followed by our framework for solving the optimization problem. We then discuss the novel design of our attacking loss and fidelity constraints, which provide more efficient criteria compared to previous methods. Finally, we introduce an advanced design to enhance adversarial image quality by latent optimization via victim-model-agnostic VAE.
Figure 3: Overview of our AtkPDM+ algorithm: Starting from the latent, $\mathbf{z}^{\operatornamewithlimits{adv}}$, of the initial adversarial image, we first decode back to pixel-domain to perform forward diffusion with both $\mathbf{x}$ and $\mathbf{x}^{\operatornamewithlimits{adv}}$ and feed them to frozen victim UNet. We then extract the feature representation of the middle block in UNet to calculate our $\mathcal{L}_{\text{attack}}$, aiming to distract the recognition of image semantics. We also calculate our $\mathcal{L}_{\text{fidelity}}$ in pixel-domain to constrain the optimization. Finally, the $\mathbf{z}^{\operatornamewithlimits{adv}}$ is being alternatively updated by loss gradients.
@@ 3.2 Overview
To achieve effective protection against diffusion-based editing, we aim to push the adversarial image away from the original clean image by disrupting the intermediate step in the reverse diffusion process. For practical real-world applications, it’s essential to ensure the adversarial image is perceptually similar to the original image. In practice, we uniformly sample the value of the forward diffusion step $t\sim[0,T]$ to generate noisy images and then perform optimization to craft the adversarial image $\mathbf{x}^{\operatornamewithlimits{adv}}$ via our attacking and fidelity losses, repeating the same process $N$ times or until convergence. Figure 2 depicts these two push-and-pull criteria during different noise levels, the successful attack is represented in the light orange line where the reverse sample moves far away from the normal edition of the image. More specifically, our method can be formulated as follows:
$\max_{\mathbf{x}^{\operatornamewithlimits{adv}}\in\mathcal{M}}\mathbb{E}_{t,\mathbf{x}_{t}|\mathbf{x},\mathbf{x}_{t}^{\operatornamewithlimits{adv}}|\mathbf{x}}\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})$ (2)
@@
Our framework, shown in Figure 3, utilizes two identical and frozen victim UNets to extract feature representations from clean and adversarial images for our attacking loss calculation and a victim-model-agnostic VAE for the latent optimization strategy.
@@ 3.3 Proposed Losses
We propose two novel losses as optimization objectives to craft an adversarial example efficiently without running through all the diffusion steps. The attacking loss is designed to distract the feature representation of the denoising UNet; The fidelity loss is a constraint to ensure the adversarial image quality. For notation simplicity, we first define the samples $\mathbf{x},\mathbf{x}^{\operatornamewithlimits{adv}}$ in different forwarded steps. Let $\mathcal{F}(\mathbf{x},t,\epsilon)=\sqrt{\bar{\alpha}_{t}}\mathbf{x}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ be the diffusion forward process. Given timestep $t$ sample from $[0,T]$, noises $\epsilon,\epsilon^{\operatornamewithlimits{adv}}$ sample from $\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote $\mathbf{x}_{t}=\mathcal{F}(\mathbf{x},t,\epsilon)$, and $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}=\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon^{\operatornamewithlimits{adv}})$.
@@
Our goal is to define effective criteria that could finally distract the reverse denoising process. PhotoGuard (Salman et al. 2023) proposed to backpropagate through all the steps of the reverse denoising process via PGD. However, this approach is prohibitively expensive, Diff-Protect (Xue et al. 2023) proposed to avoid the massive cost by leveraging Score Distillation (Poole et al. 2023) in optimization. Nevertheless, Diff-Protect relies heavily on gradients of attacking encoder of an LDM as stated in their results. In PDM, we don’t have such an encoder to attack; however, we find that the denoising UNet has a similar structure to encoder-decoder models, and some previous works (Lin and Yang 2024; Li et al. 2023) characterize this property to accelerate and enhance the generation. From our observations of the feature roles in denoising UNets, we hypothesize that distracting specific inherent feature representation in UNet blocks could lead to effectively crafting an adversarial image. In practice, we first extract the feature representations of forwarded images $\mathbf{x}_{t}$ and $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}$ in frozen UNet blocks of timestep $t$. Then, we adopt 2-Wasserstein distance (Arjovsky, Chintala, and Bottou 2017) to measure the discrepancy in the UNet feature space. The reason for choosing the 2-Wasserstein distance is that it better captures the distributional discrepancy via Optimal Transport Theory (Chen, Georgiou, and Tannenbaum 2018). Note that we aim to maximize the distance between $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}$ and $\mathbf{x}_{t}$ in the UNet feature space to distract the denoising process. Formally, the attacking loss $\mathcal{L}_{\text{attack}}$ is defined as:
$\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}^{\operatornamewithlimits{adv}}_{t})=\mathcal{W}_{2}\left(\mathcal{U}^{\text{(mid)}}_{\theta}(\mathbf{x}_{t}),\mathcal{U}^{\text{(mid)}}_{\theta}(\mathbf{x}^{\operatornamewithlimits{adv}}_{t})\right).$ (3)
Assuming the feature distributions approximate normal distributions expressed by mean $\mu_{t}$ and $\mu_{t}^{\operatornamewithlimits{adv}}$, and non-singular covariance matrices $\Sigma_{t}$ and $\Sigma_{t}^{\operatornamewithlimits{adv}}$. The calculation of the 2-Wasserstein distance between two normal distributions is viable through the closed-form solution (Dowson and Landau 1982; Olkin and Pukelsheim 1982; Chen, Georgiou, and Tannenbaum 2018):
@@
To control the attack budget for adversarial image quality, we design a constraint function that utilizes the feature extractor from a pretrained classifier to calculate the fidelity loss. In our case, we sum up the 2-Wasserstein feature losses of $L$ different layers. Specifically, we define $\mathcal{L}_{\text{fidelity}}$ as: